OpenStudy (anonymous):

Is the equation true, false, or open? Explain for more understanding, please. 9p + 8 = 10p + 7

OpenStudy (anonymous):

true for p>0 plug in p=1 17=17 false for p<0 p=-1 -9+8=-10+7 -1=-3

OpenStudy (anonymous):

So the expressions are the same for all of the values of variables?